OpenStudy (anonymous):

What are the intercepts of the line ? A. x-intercept: –3; y-intercept: 2 B. x-intercept: 3; y-intercept: –2 C. x-intercept: –2; y-intercept: 3 D. x-intercept: 2: y-intercept: –3 PLZ HELP FAST

OpenStudy (anonymous):

What line?

OpenStudy (anonymous):

6x - 9y =18

OpenStudy (anonymous):

x = 3 y = -2 so answer is B To answer this question just make x 0 and solve for y and vice versa

OpenStudy (anonymous):

sry i got d do u know y i got D?

OpenStudy (anonymous):

Was the answer D? I can't tell you how you got D if you don't show any working. X intercept is found when Y=0 so 6x - 9(0)=18 x = 18/6 x= 3 and solve for y as well.

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

got a 100% YESH

OpenStudy (anonymous):

lol Good work! Make sure you know how to do it though. When you are looking for intercepts just make x =0 to find the y intercept